OpenStudy (anonymous):

solve equation by factoring x^2+7x-44=0

OpenStudy (anonymous):

(x-4)(x+11)=0 x=4 x=-11

OpenStudy (anonymous):

\[x^2+7x-44=0\] (x+11)(x-4)=0 x=-11 0r x=4

OpenStudy (anonymous):

(x-4) (x+11) = 0